# Unraveling the Spin Coulomb Drag Effect and Its Impact on Spin Transport in the Three-Dimensional Electron Gas

### v1

@@ III.1 Ultranonlocality in the Spin Channel
A key manifestation of many-body effects in the spin response of the 3D UEG caused by the spin Coulomb drag is the ultranonlocal behavior of the dynamic XC kernel $K_{\rm XC}^{-}$. This ultranonlocality, often referred to as the ‘ultranonlocality problem’ in time-dependent spin-density-functional theory Qian et al. (2003), is characterized by a divergence of $K_{\rm XC}^{-}$ in the small-$q$ limit at finite frequencies.
To investigate this ultranonlocal behavior, we compute $K_{\rm XC}^{-}$ through the spin-spin response function $\chi_{S}$ and the non-interacting response function $\chi_{0}$ using
$K_{\rm XC}^{-}=\frac{1}{\chi_{0}}-\frac{1}{\chi_{S}}.$ (4)
Figure 3(a) shows the dimensionless $K_{\rm XC}^{-}$ as a function of $\omega_{n}$ for different momenta at $r_{s}=1.0$. Apart from the static case ($\omega_{n}$ = 0), $K_{\rm XC}^{-}$ increases rapidly as $q$ approaches zero for a given frequency, clearly demonstrating the ultranonlocal behavior.
@@
$K_{\rm XC}^{+}=\frac{1}{\chi_{0}}-\frac{1}{\chi_{nn}},$ (5)
@@
Figure 4: The reduced XC kernel $\tilde{K}_{\rm XC}^{-}$ excluding the $\mathcal{O}(\tilde{q}^{2})$ term versus $1/\tilde{q}^{2}$ ($\tilde{q}=q/k_{\rm F}$) with various frequency $\omega_{n}$ for (a) $r_{s}=0.5$, (b) $r_{s}=1.0$ and (c) $r_{s}=2.0$. The slope of the solid lines $a$ and the parameter $c$ is derived by fitting Eq. (8) for each frequency. It implies that there is a $1/q^{2}$-divergence of the spin XC kernel, which demonstrate the ultranonlocal behaviors in the spin response.
Importantly, our analysis of the singular behaviors of the XC kernel, as expressed in Eq. (2), can be carried out in the Matsubara frequency representation ${\omega_{n}}$ at the effective zero temperature. This is made possible by the application of the Wick rotation, a technique that performs an analytical continuation from real frequencies to imaginary frequencies via the transformation $\omega+i0^{+}\to i\omega_{n}$. The Wick rotation is a powerful tool in quantum field theory and many-body physics, allowing for the calculation of real-time quantities using imaginary-time formalisms Altland and Simons (2010).
The applicability of the Wick rotation in our case is justified by the fact that the XC kernel is analytic in the upper half of the complex frequency plane Giuliani and Vignale (2005). This analyticity property ensures that the imaginary-frequency representation contains the same information as the real-frequency one. Moreover, the Matsubara formalism is particularly advantageous for numerical calculations, as it avoids the singularities and branch cuts that may appear in the real-frequency domain Bruus and Flensberg (2004).
@@
$\tilde{K}_{{\rm XC}}^{-}=\frac{a}{\tilde{q}^{2}}+b+c\tilde{q}^{2},$ (8)
where $a(\omega_{n})$, $b(\omega_{n})$, and $c(\omega_{n})$ are frequency-dependent fitting parameters.
@@
Figure 5: The reduced XC kernel versus $\omega_{n}^{2}/q^{2}$ for various $r_{s}$: (a) 0.5, (b) 1.0, (c) 2.0. The slope of the solid lines represent the parameter $a_{2}$, which implies the spin mass renormalization term is dominant at low temperature while the spin diffusion is vanishing.
@@ III.2 Spin Mass Enhancement and Spin Diffusion Phenomenon
The spin mass enhancement factor $m_{s}/m$ quantifies the renormalization of the electron mass due to many-body effects in the spin channel. Our numerical analysis yields highly precise values for the spin mass enhancement factor $m_{s}/m$: $1.0072(8)$ for $r_{s}=0.5$, $1.0217(8)$ for $r_{s}=1.0$, and $1.047(10)$ for $r_{s}=2.0$. These results represent a significant advancement in the quantitative understanding of many-body effects in the spin channel. Notably, our findings for $r_{s}=1.0$ and $2.0$ are in consistent with the theoretical predictions of $m_{s}/m=1.02$ and $1.06$, respectively, as reported in Ref. Qian et al. (2004), but with substantially improved precision. The enhanced precision achieved in our calculations enables a more accurate exploration of spin-related phenomena in interacting electron systems and facilitates more stringent comparisons between theoretical predictions and experimental observations. Furthermore, our results reveal a clear trend of increasing spin mass enhancement with decreasing electron density (increasing $r_{s}$), suggesting that many-body effects become more pronounced in the strongly correlated regime.
The relaxation time of the spin diffusion $\tau_{sd}$ characterizes the decay of spin currents due to electron-electron interactions. In the low-temperature limit, $\tau_{sd}$ is expected to diverge in the 3D UEG, as the phase space for electron-electron scattering vanishes D’Amico and Vignale (2000); Giuliani and Vignale (2005). Our numerical results confirm this behavior, as the ratio of the inverse Fermi energy to $\tau_{sd}$, is found to be of the order of $10^{-3}$ for $r_{s}=0.5$ and $1.0$. For $r_{s}=2.0$, the uncertainty in $a_{1}$ exceeds its value, making a reliable determination of $\tau_{sd}$ challenging. Nevertheless, the overall trend suggests that $1/\tau_{sd}$ approaches zero as the temperature tends to zero, consistent with the theoretical expectation of vanishing spin diffusion in the 3D UEG at zero temperature.
@@
In summary, our numerical analysis of the frequency-dependent fitting parameter $a(\omega_{n})$ provides strong evidence for the presence of spin mass enhancement and the suppression of spin diffusion in the 3D UEG at low temperatures. The precise determination of the spin mass enhancement factor and the confirmation of the vanishing spin diffusion in the zero-temperature limit showcase the power of the VDMC approach in capturing subtle many-body effects in the spin channel. These findings contribute to a deeper understanding of spin transport phenomena in interacting electron systems and serve as a benchmark for future theoretical and computational studies.
@@ IV Discussions
In this work, we have employed the Variational Diagrammatic Monte Carlo (VDMC) approach to investigate the spin response of the three-dimensional uniform electron gas (3D UEG) at low temperatures. Our study has focused on the spin Coulomb drag effect through the ultranonlocality behaviors of the spin-resolved exchange-correlation (XC) kernel and its connection to fundamental properties of spin transport, namely the spin mass enhancement and the spin diffusion phenomenon.
Through extensive VDMC simulations, we have computed the spin-spin response function and the associated XC kernel for a wide range of electron densities, momenta, and Matsubara frequencies. Our results clearly demonstrate the presence of a $1/q^{2}$ divergence in the spin XC kernel at finite frequencies, confirming the spin Coulomb drag effect predicted by theoretical studies. Remarkably, this singularity is absent in the charge channel, highlighting the unique nature of many-body effects in the spin response.
By fitting our numerical data to an analytical ansatz for the XC kernel, we have extracted the frequency-dependent coefficient of the $1/q^{2}$ term, which encodes information about the spin mass enhancement and the dissipation of the spin current. Our analysis has yielded highly precise values for the spin mass enhancement factor, surpassing the accuracy of previous theoretical predictions. Interestingly, we have observed a trend of increasing spin mass enhancement with decreasing electron density, indicating the growing importance of many-body effects in the strongly correlated regime.
Furthermore, our study has provided compelling evidence for the suppression of spin diffusion in the 3D UEG at low temperatures. The extracted relaxation times of the spin current are found to be several orders of magnitude larger than the Fermi time, suggesting that spin currents can persist for extended durations in this system. This finding is consistent with the theoretical expectation of vanishing spin diffusion in the zero-temperature limit.
The results presented in this work represent a significant advancement in the quantitative understanding of spin-related phenomena in interacting electron systems. The unprecedented precision achieved in our calculations opens up new possibilities for probing subtle many-body effects and facilitates more stringent comparisons between theory and experiment. Moreover, our study highlights the power and versatility of the VDMC approach in tackling challenging problems in the field of spin transport.
Looking ahead, the insights gained from this work can guide the development of more accurate density functional approximations for spin-dependent phenomena, with potential applications in spintronics and quantum technologies. The extension of our methodology to more complex systems, such as multicomponent electron gases and realistic materials, presents exciting avenues for future research.
In conclusion, our comprehensive investigation of the spin response in the 3D UEG has unveiled the intricate interplay between ultranonlocality, spin mass enhancement, and spin Coulomb drag. The precise numerical results and the rigorous theoretical analysis presented here establish a new benchmark for the study of spin transport in interacting electron systems and pave the way for future discoveries in this fascinating field.
@@ Acknowledgements
Z.L, P.H. and Y.D. were supported by the supported by the National Natural Science Foundation of China (under Grant No. 12275263), the Innovation Program for Quantum Science and Technology (under grant No. 2021ZD0301900), the Natural Science Foundation of Fujian Province of China (under Grant No. 2023J02032). K.C. was supported by National Natural Science Foundation of China under Grants No. 12047503.

### v4

@@ III.1 Ultranonlocality in the Spin Channel
A key manifestation of many-body effects in the spin response of the 3D UEG caused by the spin Coulomb drag is the ultranonlocal behavior of the dynamic XC kernel $K_{\rm XC}^{-}$. This ultranonlocality, often referred to as the ‘ultranonlocality problem’ in time-dependent spin-density-functional theory [16], is characterized by a divergence of $K_{\rm XC}^{-}$ in the small-$q$ limit at finite frequencies.
To investigate this ultranonlocal behavior, we compute $K_{\rm XC}^{-}$ through the dynamic spin susceptibility $\chi_{S}$ and the Lindhard function $\chi_{0}$ using Eq. (1). Figure 3(a) shows the dimensionless $K_{\rm XC}^{-}$ as a function of $\omega_{n}$ for different momenta at $r_{s}=1.0$. Apart from the static case ($\omega_{n}$ = 0), $K_{\rm XC}^{-}$ increases rapidly as $q$ approaches zero for a given frequency, clearly demonstrating the ultranonlocal behavior.
@@
$K_{\rm XC}^{+}={\chi_{0}}^{-1}(\mathbb{q},\omega)-\chi^{-1}_{nn}(\mathbb{q},\omega),$ (5)
@@
Figure 4: The reduced XC kernel $\tilde{K}_{\rm XC}^{-}$ excluding the $\mathcal{O}(\tilde{q}^{2})$ term versus $1/\tilde{q}^{2}$ ($\tilde{q}=q/k_{\rm F}$) with various frequency $\omega_{n}$ for (a) $r_{s}=0.5$, (b) $r_{s}=1.0$ and (c) $r_{s}=2.0$. The slope of the solid lines $a$ is derived by fitting Eq. (8) for each frequency. It implies that there is a $1/q^{2}$-divergence of the spin XC kernel, which demonstrate the ultranonlocal behaviors in the spin response.
Importantly, our analysis of the singular behaviors of the XC kernel, as expressed in Eq. (3), can be carried out in the Matsubara frequency representation ${\omega_{n}}$ at the effective zero temperature. This is made possible by the application of the Wick rotation, a technique that performs an analytical continuation from real frequencies to imaginary frequencies via the transformation $\omega+i0^{+}\to i\omega_{n}$. The Wick rotation is a powerful tool in quantum field theory and many-body physics, allowing for the calculation of real-time quantities using imaginary-time formalisms [52].
The applicability of the Wick rotation in our case is justified by the fact that the XC kernel is analytic in the upper half of the complex frequency plane [53]. This analyticity property ensures that the imaginary-frequency representation contains the same information as the real-frequency one. Moreover, the Matsubara formalism is particularly advantageous for numerical calculations, as it avoids the singularities and branch cuts that may appear in the real-frequency domain [54].
@@
$\tilde{K}_{{\rm XC}}^{-}(\tilde{q};\omega_{n})=\frac{a(\omega_{n})}{\tilde{q}^{2}}+b(\omega_{n}),$ (8)
where $a(\omega_{n})$, $b(\omega_{n})$ are frequency-dependent fitting parameters. Given that our data falls within the small-$q$ range ($0.1\leq\tilde{q}\leq 0.5$), we omit terms of $q^{2}$ and higher orders to ensure a stable fit.
@@
Figure 5: The reduced XC kernel versus $\tilde{\omega}^{2}/\tilde{q}^{2}$ for various $r_{s}$: (a) 0.5, (b) 1.0, (c) 2.0. The slope of the dashed lines represent the parameter $a_{2}$ and the shaded area represents its error, which implies the spin mass renormalization term is dominant at low temperature while the spin diffusion is vanishing.
@@ III.2 Spin Mass Enhancement and Spin Diffusion Phenomenon
The spin mass enhancement factor $m_{s}/m$ quantifies the renormalization of the electron mass due to many-body effects in the spin channel. Our numerical analysis yields highly precise values for the spin mass enhancement factor $m_{s}/m$: 1.007 7(8)1.007781.007\,7(8)1.007 7 ( 8 ) for $r_{s}=0.5$, 1.019 3(3)1.019331.019\,3(3)1.019 3 ( 3 ) for $r_{s}=1.0$, and $1.04(2)$ for $r_{s}=2.0$. These results corroborate previous theoretical predictions of $m_{s}/m=1.02$ and $1.06$ for $r_{s}=1.0$ and $2.0$ respectively [20], which were obtained from the calculation of the dynamic spin XC kernel through a decoupling approximation of the four-point density function combined with diffusion Monte Carlo-calculated static local field factors [31, 22]. Our calculations achieve substantially higher precision, enabling more rigorous tests of theoretical models against experimental measurements. The systematic increase in spin mass enhancement with $r_{s}$ indicates strengthening many-body effects as the electron system becomes more strongly correlated.
The relaxation time of the spin diffusion $\tau_{sd}$ characterizes the decay of spin currents due to electron-electron interactions. In the low-temperature limit, $\tau_{sd}$ is expected to diverge in the 3D UEG, as the phase space for electron-electron scattering vanishes [1, 53]. Our numerical results confirm this behavior, as the ratio of the inverse Fermi energy to $\tau_{sd}$, is found to be of the order of $10^{-4}$ for $r_{s}=0.5$ and $1.0$. For $r_{s}=2.0$, the uncertainty in $a_{1}$ exceeds its value, making a reliable determination of $\tau_{sd}$ challenging. Nevertheless, the overall trend suggests that $1/\tau_{sd}$ approaches zero as the temperature tends to zero, consistent with the theoretical expectation of vanishing spin diffusion in the 3D UEG at zero temperature.
@@
Through systematic analysis of the frequency-dependent parameter $a(\omega_{n})$, our VDMC calculations quantify the spin mass enhancement and spin diffusion suppression in the three-dimensional uniform electron gas at low temperatures. The precise determination of these quantities demonstrates the capability of VDMC to resolve subtle many-body effects in spin transport, providing quantitative benchmarks for future theoretical and computational studies.
@@ IV Discussions
In this work, we have employed the VDMC approach to investigate the spin susceptibility of the 3D UEG at low temperatures. Our study has focused on the spin Coulomb drag effect through the ultranonlocality behaviors of the spin-resolved XC kernel and its connection to fundamental properties of spin transport, namely the spin mass enhancement and the spin diffusion phenomenon.
Through extensive VDMC simulations, we have computed the dynamic spin susceptibility and the associated XC kernel for various electron densities, momenta, and Matsubara frequencies. Our results clearly demonstrate the presence of a $1/q^{2}$ divergence in the spin XC kernel at finite frequencies, confirming the spin Coulomb drag effect predicted by theoretical studies. Remarkably, this singularity is absent in the charge channel, highlighting the unique nature of many-body effects in the spin response.
By fitting our numerical data to an analytical ansatz for the XC kernel, we have extracted the frequency-dependent coefficient of the $1/q^{2}$ term, which encodes information about the spin mass enhancement and the dissipation of the spin current. Our analysis yields precise values for the spin mass enhancement factor, aligning closely with previous theoretical predictions and significantly improving upon their accuracy. We observe an increasing trend in spin mass enhancement with decreasing electron density, highlighting the growing importance of many-body effects in the strongly correlated regime. Furthermore, our study presents numerical evidence for the suppression of spin diffusion in the 3D UEG at low temperatures. The extracted relaxation times of the spin current are found to be several orders of magnitude larger than the inverse Fermi energy, suggesting that spin currents can persist for extended durations in this system. This finding is consistent with the theoretical expectation of vanishing spin diffusion in the zero-temperature limit [53].
Looking ahead, the insights gained from this work may guide the development of more accurate density functional approximations for spin-dependent phenomena, with potential applications in spintronics and quantum technologies. Future extensions of our methodology to more complex systems, such as multicomponent electron gases, cold atomic gases, and realistic materials, hold promise for further advancing the field.
@@ Acknowledgements
K. C. was supported by the National Key Research and Development Program of China, Grant No. 2024YFA1408604, and the National Natural Science Foundation of China under Grants No. 12047503 and No. 12447103. Z.L, P.H., and Y.D. were supported by the National Natural Science Foundation of China (under Grant No. 12275263), the Innovation Program for Quantum Science and Technology (under grant No. 2021ZD0301900), the Natural Science Foundation of Fujian Province of China (under Grant No. 2023J02032).